OpenStudy (anonymous):

The mass of a lift is 600kg. The max tensile force that the cable supporting the lift can withstand is 7kN. calculate the max upward acceleration allowed.

OpenStudy (anonymous):

11.666666666666666666666666666667 m/sec2

OpenStudy (anonymous):

can you please show your calculation?

OpenStudy (anonymous):

what formulae did you use?

OpenStudy (anonymous):

the answer is 1.867 m/s2. just dont know how to calculate it.

OpenStudy (anonymous):

You sure? I get .86. Its a tension problem so it should be in equilibrium. I am getting .86 m/s2

OpenStudy (anonymous):

are'nt you forgetting about gravity? 10 m/s2

OpenStudy (anonymous):

1.867 is the answer in my book so should be right

OpenStudy (anonymous):

T= W + ma= mg +ma

OpenStudy (anonymous):

T= 7000N = 600 (10 +a) 70/6= 10 +a 70/6 -10 =a 11.67 -10= a 1.67= a

OpenStudy (anonymous):

They used 9.8 not 10, so you do get the 1.87 m/s2

OpenStudy (anonymous):

well done! thanks.